For each molecular substance (a) H3PO2 and (b) C6H5NH2, calculate its molecular mass and write a conversion factor that converts between mass in grams and moles of the substance.
Answer #1

1) H3PO2

Molar mass of H = 1 g/mol

Molar mass of P = 31 g/mol

Molar mass of O = 16 g/mol

So, Molecular mass of H3PO2 = 3*1 + 31 + 2*16 = 66 g/mol .... Answer

2) C6H5NH2

Molar mass of C = 12 g/mol

Molar mass of N = 14 g/mol

So, molar mass of C6H5NH2 = 12*6 + 5*1 + 14 + 1*2 = 93 g/mol .... Answer

Relationship between mass in grams and moles

Moles of compound(n) = mass of compound in grams(m) / molar mass of compound(M)

i.e. n = m/M